OpenStudy (anonymous):

I have to determine the end behavior of the graph of each polynomial function. I'm just really confused on how to graph variables with exponents..

OpenStudy (anonymous):

\[y=x ^{2}-2x+3\] \[y=\frac{ 1 }{ 2 }x^{4}+5x ^{2}-\frac{ 1 }{ 2 }\]

OpenStudy (whpalmer4):

Same way you would graph variables without exponents. You plug in the number for \(x\) and find the corresponding value for \(y\). In your first equation, \(y = x^2-2x+3\) \(x = 0, ~y = (0)^2-2(0)+3 = 0-0+3\) \(x=1,~y = (1)^2 - 2(1)+3 = 1-2+3 = 2\) \(x=2,~y = (2)^2-2(2)+3) = 4-4+3=3\) etc. You should also plot some points with negative values of \(x\).

OpenStudy (whpalmer4):

But to find end behavior, you don't really need to graph anything — just look at the term with the largest exponent and decide what it does at very positive and very negative values. For example, if our function happened to be \(y = 3x^3 + 2x^2 -5x + 4\), at large values of \(x\) (both positive and negative) the \(3x^3\) term is going to dominate the result. Consider \(x=100\): the respective terms are \(3(100)^3, 2(100)^2,-5(100),4\) or \(3000000, 20000,-500,4\) As \(x\) gets closer to positive or negative infinity, the value will be almost entirely that of the \(3x^3\) term. So, for my example (which is NOT your problem), the end behavior is down and to the left to negative infinity at negative values of \(x\), and up and to the right to positive infinity at positive values of \(x\), because that is the end behavior of \(3x^3\). Plug in a big positive number, you get a big positive number in return. Plug in a big negative number, you get a big negative number in return. Any questions?